OpenStudy (anonymous):

lim (3)/(y+4) y --> -infinity

jimthompson5910 (jim_thompson5910):

y is approaching negative infinity so y+4 is also approaching negative infinity

jimthompson5910 (jim_thompson5910):

the denominator is some number that keeps getting larger and larger in the negative direction

OpenStudy (anonymous):

\[\lim_{y \rightarrow - \infty }\frac{ 3 }{ y+4 }=\lim_{y \rightarrow - \infty }\frac{ \frac{ 3 }{ y } }{ 1+\frac{ 4 }{ y }}=\frac{ 0 }{ 1+0 }=0\]

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

yw